OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A. y-1=6(x+5) B. y+5=6(x-1) C. y+1=6(x-5) D. y-5=6(x+1)

OpenStudy (anonymous):

y-y1=m(x-x1) where m=slope (x1,y1) is your point igven to you.

OpenStudy (anonymous):

Point-slope form: y − b = m(x − a) for a point (a, b) and slope m Slope-intercept form: y = mx + b, for slope m and y-intercept b

OpenStudy (cwrw238):

general form is y-y1 = m(x-x1) where m = slope and (x1,y1) is a point on the line

OpenStudy (cwrw238):

so you have m = 6 x1 = 5 and y1 = -1 so now you should be able to pick the correct choice